OpenStudy (gigi1103):

y−2=−2(x+4) how do I plot this / can you check my answers

OpenStudy (shadowlegendx):

Distribute -2 to x and 4 Then add 2 to both sides

OpenStudy (shadowlegendx):

\[y - 2 = -2(x + 4)\] \[y - 2 = -2x - 8\] \[y = -2x - 6\]

OpenStudy (gigi1103):

What is the equation, in point-slope form, for a line that goes through (8, −4) and has a slope of −5/6 ? y−4=−5/6(x+8)? What is the equation of a line, in point-slope form, that passes through (−2, −6) and has a slope of 1/3? y−2=13(x−6)? What is the equation in point-slope form of a line that passes through the points (−4, −1) and (5, 7)? What is the equation in point-slope form of a line that passes through the points 5, −3) and (−2, 9) ? y−3=−127(x+5)

OpenStudy (gigi1103):

so is the 6 negative also?

OpenStudy (shadowlegendx):

To plot, go to -6 on the y axis, plot a point. Next, from that first point, go down 2 levels on the y axis, then go to the right on the x axis. Then plot your second point. For your third point. From your first point, go up 2 levels on the y axis, then 1 level to the left on the x axis

OpenStudy (gigi1103):

thanks!

OpenStudy (shadowlegendx):



OpenStudy (gigi1103):

thank you that really helped can you help check my answers?

OpenStudy (shadowlegendx):

I have to go, but just follow the steps that I showed you, and you should be fine.

OpenStudy (gigi1103):

ok thanks!